# E-values, Multiple Testing and Beyond

### v1

@@ Abstract
We discover a connection between the Benjamini-Hochberg (BH) procedure and the recently proposed e-BH procedure Wang and Ramdas (2022) with a suitably defined set of e-values. This insight extends to a generalized version of the BH procedure and the model-free multiple testing procedure in Barber and Candès (2015) (BC) with a general form of rejection rules. The connection provides an effective way of developing new multiple testing procedures by aggregating or assembling e-values resulting from the BH and BC procedures and their use in different subsets of the data. In particular, we propose new multiple testing methodologies in three applications, including a hybrid approach that integrates the BH and BC procedures, a multiple testing procedure aimed at ensuring a new notion of fairness by controlling both the group-wise and overall false discovery rates (FDR), and a structure adaptive multiple testing procedure that can incorporate external covariate information to boost detection power. One notable feature of the proposed methods is that we use a data-dependent approach for assigning weights to e-values, significantly enhancing the efficiency of the resulting e-BH procedure. The construction of the weights is non-trivial and is motivated by the leave-one-out analysis for the BH and BC procedures. In theory, we prove that the proposed e-BH procedures with data-dependent weights in the three applications ensure finite sample FDR control. Furthermore, we demonstrate the efficiency of the proposed methods through numerical studies in the three applications.
Keywords: Benjamini-Hochberg procedure, Cross-fitting, E-values, False discovery rate, Leave-one-out analysis, Multiple testing
@@ 1 Introduction
When working with high-dimensional data in modern scientific fields, the problem of multiple testing often arises when we explore a vast number of hypotheses with the goal of detecting signals while also controlling some error measures, such as the false discovery rate (FDR). The Benjamini-Hochberg (BH) procedure (Benjamini and Hochberg, 1995) is perhaps the most widely used FDR-controlling procedure that rejects a hypothesis whenever its p-value is less than or equal to an adaptive rejection threshold determined by the whole set of p-values. Barber and Candès (2015) proposed a model-free FDR-controlling (BC) procedure that estimates the number of false rejections by leveraging the symmetry of p-values under the null and compares each p-value with an adaptive threshold. More recently, there is a growing literature on utilizing e-values for statistical inference under different contexts, see, e.g., (Grünwald et al., 2020; Shafer, 2021; Vovk and Wang, 2021; Xu et al., 2021; Ignatiadis et al., 2022; Dunn et al., 2023; Xu and Ramdas, 2023). In particular, Wang and Ramdas (2022) proposed a multiple testing procedure (named e-BH procedure) by applying the BH procedure to e-values, which was shown to control the FDR even when the e-values exhibit arbitrary dependence.
In this work, we establish a connection between the BH and e-BH procedures with a suitably defined set of e-values, proving that they yield identical rejection sets. This connection extends to a generalized version of the BH and BC procedures, which can have a more general form for the rejection rule. Based on these findings, we propose new multiple testing procedures by aggregating e-values from the BH and BC procedures or assembling e-values from different subsets of the data. Specifically, we consider three concrete applications, which we will illustrate in detail below.
@@
Our first application concerns the development of a robust and efficient multiple testing procedure by leveraging the strength of the BH and BC procedures. The BH and BC procedures employ different strategies to estimate the number of false rejections, leading to distinct performances depending on the underlying signal density and strength. Empirical results in the literature suggest that neither one dominates the other. For instance, the BH procedure is usually more powerful when the underlying signal is sparse, while the BC procedure excels in the case of dense signal (Arias-Castro and Chen, 2017). It is often impossible to know in advance which method will perform better in real-world applications. Reporting results from the better-performing method can lead to an inflated FDR and is considered as a type of data snooping. To develop a more efficient procedure with guaranteed FDR control, we propose a hybrid method that utilizes e-values as a bridge to integrate the testing results from the BH and BC procedures. Specifically, we construct a set of e-values by properly weighting the e-values from both the BH and BC procedures. We then use the constructed e-values as the input for the e-BH procedure. We show that the resulting hybrid procedure controls the FDR and can significantly improve the worse-performing method in finite sample.
Our second application is concerned with a setting where one has to make (high-stakes) decisions by testing $n$ hypotheses that are partitioned into $G$ groups according to some (protected) attributes. Let us consider the example of a bank’s loan approval process, where customers are grouped by gender or race. The null hypothesis in this case is that the loan should be approved for the customer. The bank faces the challenge of controlling the overall FDR, which means that the bank should not falsely decline too many customers’ applications, as it would reduce the loan volume and decrease income. Additionally, the bank also needs to control the group-wise FDR to ensure that certain racial or gender groups are not unfairly denied loans. If the group-wise FDR inflates, customers who believe they are unfairly denied loans may take their business elsewhere. Therefore, the bank aims to maintain control over the FDR for each group while also controlling the overall FDR across all hypotheses. Applying an FDR-controlling procedure to the whole set of hypotheses, ignoring the group structure will fail to control the FDR for individual groups. On the other hand, performing multiple testing for each group separately at level $\alpha$ does not ensure that the overall FDR is controlled at the same level $\alpha$. A naive procedure to achieve error control at both the group and overall levels is to conduct multiple testing for each group separately at the reduced level $\alpha/G$, which turns out to be too conservative, as seen in our (unreported) numerical studies. To address this challenge, we propose a fair multiple testing procedure that uses e-values as a bridge to assemble the results from different groups, controlling the FDR within each group and the overall FDR simultaneously. Specifically, we implement the BC procedure for each group, followed by constructing e-values by assembling e-values from different groups with proper weights, and using the constructed e-values as the input for the e-BH procedure. We show that the resulting procedure simultaneously controls the FDR within each group and the overall FDR in finite sample.
@@
As the final application, we consider the problem of multiple testing with external structural information in the form of covariates, which has received much recent attention, as leveraging auxiliary information can enhance the power and interpretability of multiple testing results in many scientific applications. For instance, in differential expression analysis of RNA-seq data, which tests for differences in the mean expression of the genes between conditions, the sum of read counts per gene across all samples could be the auxiliary data since it is informative of the statistical power. In differential abundance analysis of microbiome sequencing data, which tests for differences in the mean abundance of the detected bacterial species between conditions, the genetic divergence among species is important auxiliary information since closely related species usually have similar physical characteristics and tend to covary with the condition of interest. A growing list of works has reflected the importance of this research direction in recent years, for instance, (Genovese et al., 2006; Hu et al., 2010; Sun et al., 2015; Ignatiadis et al., 2016; Boca and Leek, 2018; Lei and Fithian, 2018; Li and Barber, 2019; Ignatiadis and Huber, 2021; Yun et al., 2022; Zhang and Chen, 2022; Zhao and Zhou, 2023). However, these existing works suffer from different limitations. For example, the local FDR-based method (Sun et al., 2015; Cao et al., 2022) lacks finite sample FDR control and can only guarantee FDR control asymptotically. The weighted BH methods (Ignatiadis and Huber, 2021; Li and Barber, 2019) lead to suboptimal power, as observed in our numerical studies. To address these drawbacks, we propose a powerful multiple testing procedure that incorporates auxiliary information with guaranteed FDR control in finite samples. We randomly split samples into several groups and use the cross-fitting approach to estimate the rejection function in each group using all samples in other groups. Then, we implement the generalized BC procedure for each group. We assemble the BC procedure induced e-values from different groups with proper weights. The assembled e-values are used as the input for e-BH procedure. We show that the proposed procedure controls the FDR at the desired level in finite sample.
@@
Our approach involves a data-dependent method for weighting e-values when aggregating them from the BH and BC procedures (the first application) or assembling them from different subsets of the data (the second and third applications). It is important to note that our weighting method differs from the “boosting factor” proposed by Wang and Ramdas (2022), which involves multiplying each e-value by a factor to boost them up before applying the e-BH procedure. In our approach, we construct weights that ensure the weighted e-values satisfying Condition (3) below. This condition is sufficient to maintain FDR control of the e-BH procedure at the desired level. Our numerical findings show that implementing the e-BH procedure with data-dependent weights improves its efficiency across all applications compared to the unweighted e-BH procedure.
@@
The rest of the paper is organized as follows. In Section 2, we provide a brief review of the BH, BC, and e-BH procedures and establish the equivalence between the BH/BC and the e-BH procedures. In Section 3, we extend this equivalence to a generalized version of the BH and BC procedures. We present new multiple testing methodologies in three different applications, including a hybrid approach that integrates the BH and BC procedures, a multiple testing procedure that controls both the group-wise and overall FDRs, and a structure adaptive multiple testing procedure, in Sections 4.1, 4.2, and 4.3, respectively. Section 5 concludes. The appendix contains additional numerical results and all proofs of the main results.
@@ 2 Preliminaries
To begin with, we provide a brief overview of the BH, BC, and e-BH procedures, and establish the equivalence between the BH (BC) procedure and the e-BH procedure with a suitably defined set of e-values. This equivalence appears to be a new finding that has not been explicitly stated in the previous literature.
We are interested in testing $n$ hypotheses $\{H_{1},\dots,H_{n}\}$ simultaneously. Let $\bm{\theta}=\{\theta_{1},\dots,\theta_{n}\}\in\{0,1\}^{n}$ indicate the underlying truth of each hypothesis, where $\theta_{i}=0$ if $H_{i}$ is under the null and $\theta_{i}=1$ otherwise. Denote by $\bm{\delta}=(\delta_{1},\dots,\delta_{n})\in\{0,1\}^{n}$ a decision rule for the $n$ hypotheses, where we reject (accept) the $i$th hypothesis if $\delta_{i}=1$ ($\delta_{i}=0$). The FDR for the decision rule $\bm{\delta}$ is defined as the expectation of the false discovery proportion (FDP), i.e.,
@@
$\mathrm{FDR}(\bm{\delta})=\mathbb{E}[\mathrm{FDP}(\bm{\delta})],\quad\mathrm{FDP}(\bm{\delta})=\frac{\sum_{i=1}^{n}(1-\theta_{i})\delta_{i}}{1\vee\sum_{i=1}^{n}\delta_{i}},$
where $a\vee b=\max\{a,b\}$. The goal of an FDR controlling procedure is to ensure that the FDR is bounded from above by a pre-specified number $\alpha\in(0,1)$.
2.1 BH procedure
The Benjamini-Hochberg (BH) procedure (Benjamini and Hochberg, 1995) is perhaps the most widely used FDR-controlling method. To describe the procedure, suppose we observe a p-value $p_{i}$ for each $H_{i}$. Sort the p-values in ascending order as $p_{(1)}\leq\dots\leq p_{(n)}$ and let $\hat{k}=\max\left\{i\colon p_{(i)}\leq(\alpha i)/n\right\}$. The BH procedure rejects all hypotheses $H_{(i)}$ with $i\leq\hat{k}$, where $H_{(i)}$ is the hypothesis associated with $p_{(i)}$. This procedure is equivalent to rejecting all $H_{i}$ with $p_{i}\leq T_{\mathrm{BH}}$, where $T_{\mathrm{BH}}$ is defined as
$T_{\mathrm{BH}}=\sup\left\{0<t\leq 1\colon\frac{nt}{1\vee R(t)}\leq\alpha\right\},$ (1)
with $R(t)=\sum_{i=1}^{n}\mathbbm{1}\{p_{i}\leq t\}$ being the number of rejections given the threshold $t$, and $\mathbbm{1}\{A\}$ denoting the indicator function associated with a set $A$. It is well known that if the null p-values are mutually independent and super-uniform, and are independent of the alternative p-values, the BH procedure at level $\alpha$ controls the FDR at the level $\alpha n_{0}/n\leq\alpha$, where $n_{0}$ is the number of hypotheses under the null.
2.2 BC procedure
In a seminal paper by (Barber and Candès, 2015), the authors proposed a model-free multiple testing procedure (BC procedure hereafter) that exploits the symmetry of the null p-values (or test statistics) to estimate the number of false rejections. More precisely, the BC procedure specifies a data-dependent threshold, denoted by $T_{\mathrm{BC}}$, which is determined as follows:
$T_{\mathrm{BC}}=\sup\left\{0<t<0.5\colon\frac{1+\sum^{n}_{i=1}\mathbbm{1}\{p_{i}\geq 1-t\}}{1\vee R(t)}\leq\alpha\right\},$ (2)
and it rejects all $H_{i}$ with $p_{i}\leq T_{\mathrm{BC}}$. The BC procedure has been shown to provide finite sample FDR control under suitable assumptions (Barber and Candès, 2015).
2.3 E-values and e-BH procedure

### v2

# A General Framework for Multiple Testing via E-value Aggregation and Data-Dependent Weighting

@@ Abstract
Motivated by recent findings in Li and Zhang (2025), which established an equivalence between certain p-value-based multiple testing procedures and the e-Benjamini-Hochberg procedure (Wang and Ramdas, 2022), we introduce a general framework for constructing novel multiple testing methods through the aggregation and combination of e-values. Specifically, we propose methodologies for three distinct scenarios: (i) assembly of e-values obtained from different subsets of data, simultaneously controlling group-wise and overall false discovery rates; (ii) aggregation of e-values derived from different procedures or the same procedure employing different test statistics; and (iii) adaptive multiple testing methods that incorporate external structural information to enhance statistical power. A notable feature of our approach is the use of data-dependent weighting of e-values, significantly improving the efficiency of the resulting e-Benjamini-Hochberg procedures. The construction of these weights is non-trivial and inspired by leave-one-out analysis, a widely utilized technique for proving false discovery rate control in p-value-based methodologies. We theoretically establish that the proposed e-Benjamini-Hochberg procedures, when equipped with data-dependent weights, guarantee finite-sample false discovery rate control across all three considered applications. Additionally, numerical studies illustrate the efficacy and advantages of the proposed methods within each application scenario.
Keywords: Cross-fitting, E-values, False discovery rate, Leave-one-out analysis, Multiple testing
@@ 1 Introduction
In modern scientific research involving high-dimensional data, multiple testing frequently arises as a fundamental challenge. This occurs when simultaneously evaluating a large number of hypotheses to identify significant signals, necessitating careful control of error rates such as the false discovery rate (FDR) to ensure statistical validity.
The Benjamini-Hochberg (BH) procedure (Benjamini and Hochberg, 1995) and the Barber-Candès (BC) procedure (Barber and Candès, 2015) are the most commonly employed methods for controlling the FDR using p-values. Recently, there has been increasing interest in employing e-values for FDR control; see, for example, Ignatiadis et al. (2024a, b); Xu and Ramdas (2024). In particular, Wang and Ramdas (2022) proposed a multiple testing approach called the e-Benjamini-Hochberg (e-BH) procedure, which applies the BH procedure directly to e-values. They demonstrated that the e-BH procedure controls the FDR even when e-values exhibit arbitrary dependence structures. Compared to p-values, which are defined through tail probabilities, e-values, defined via expectation, offer greater flexibility for combining multiple e-values to obtain new e-values (Vovk and Wang, 2021). For comprehensive reviews of methods based on e-values, we refer readers to Ramdas and Wang (2024).
In recent work, Li and Zhang (2025) introduced a unified framework for multiple testing procedures based on p-values, which includes the BH and BC procedures as special cases. The authors established the equivalence between these p-value-based methods and the e-BH procedure when appropriate sets of e-values are utilized. Here, equivalence means that these methods yield identical rejection sets. Motivated by these findings, we propose new multiple testing procedures that aggregate e-values derived from different methods or the same method with different test statistics, or combine e-values obtained from different subsets of data. Specifically, we explore three concrete scenarios, which are detailed in the subsequent sections.
In our first scenario, we consider the setting with $L$ sets of e-values derived from $L$ distinct datasets. Our goal is to aggregate these e-values into a single vector to incorporate all group-level information. We propose a procedure designed to control both the overall FDR and the group-wise FDR simultaneously.
A key application involves high-stakes decision-making, where we test $n$ hypotheses partitioned into $G$ groups according to certain protected attributes. In a loan-approval system, for example, customers may be grouped by gender or race; the null hypothesis states that a given loan should be approved. Management must control both the overall FDR—so that too few loans are rejected—and the group-wise FDR—so that no group is unfairly treated. Conventional solutions fail: applying the FDR-controlling method to all $n$ tests fails to control FDR for certain groups; testing each group at level $\alpha$ fails to control overall FDR; and a Bonferroni adjustment to $\alpha/G$ is too conservative. To address this challenge, we propose a multiple-testing procedure that uses e-values as a bridge to combine the results from different groups, controlling the FDR within each group and the overall FDR simultaneously. Specifically, we apply the BC procedure to each group, then assemble the resulting e-values with appropriate weights to form a unified e-value vector, which we pass to the e-BH procedure. We show that the resulting method simultaneously controls the FDR within each group and the overall FDR in finite samples.
In our second scenario, we consider a setting with $L$ distinct sets of e-values, which we aim to aggregate into a single e-value vector to incorporate all available information. By effectively combining e-values from diverse sources, the proposed approach enables the integration of multiple results while rigorously controlling the overall FDR. This general scenario has several specific applications.
The first specific application is the development of a robust and efficient knockoff method capable of accommodating various underlying relationships between the response variable and predictors. In the knockoff methodology (Barber and Candès, 2015), several test statistics can be employed: Lasso-based methods excel for (near-)linear models, whereas random forest-based methods suit nonlinear settings (Candes et al., 2018). Because the true form of the dependence is rarely known, we develop a robust and efficient knockoff procedure that can leverage the strengths of different test statistics simultaneously. Specifically, we construct a unified set of e-values by aggregating the e-values derived from knockoff methods based on various test statistics. Subsequently, we input these aggregated e-values into the e-BH procedure. We demonstrate that the resulting hybrid knockoff procedure effectively controls the FDR and maintains robust performance regardless of the true underlying relationship between the response variable and predictors.
@@
Our second application involves developing a robust and efficient multiple-testing procedure by combining the strengths of the BH and BC procedures. The BH and BC procedures use different strategies for estimating the number of false rejections, leading to distinct performance characteristics depending on signal density and strength. In real-world applications, reporting results from the better-performing method can lead to inflated FDR and is considered a form of data snooping. To address this issue, we propose a hybrid approach that employs e-values as a bridge to integrate results from both the BH and BC procedures. To this end, we construct a unified set of e-values by suitably weighting the e-values derived from the BH and BC procedures. These combined e-values are then input into the e-BH procedure. We show that the resulting hybrid procedure maintains rigorous FDR control and can significantly improve performance relative to the weaker individual method in finite-sample settings.
@@
In the final scenario, we consider the problem of multiple testing with external structural information in the form of covariates, which has received significant recent attention, as leveraging auxiliary information can enhance the power and interpretability of multiple-testing results in many scientific applications. Typical covariates include (i) total read counts in RNA-seq, which modulate gene-level power, and (ii) phylogenetic distances in microbiome studies, where related species share abundance patterns. A growing list of works has reflected the importance of this research direction in recent years—for instance, Hu et al. (2010); Ignatiadis et al. (2016); Lei and Fithian (2018); Li and Barber (2019); Ignatiadis and Huber (2021); Zhang and Chen (2022); Zhao and Zhou (2024). However, these existing works suffer from various limitations. For example, the local FDR-based methods (Sun et al., 2015; Cao et al., 2022) lack finite-sample FDR control and only guarantee FDR control asymptotically. The weighted BH methods (Ignatiadis and Huber, 2021; Li and Barber, 2019) lead to suboptimal power, as observed in our numerical studies. To address these drawbacks, we propose a powerful multiple-testing procedure that incorporates auxiliary information while guaranteeing FDR control in finite samples. Specifically, we randomly split the data into several disjoint groups and use a cross-fitting approach (Ignatiadis and Huber, 2021) to estimate the rejection function for each group using all samples in the other groups. We then apply the flexible BC procedure (Li and Zhang, 2025) within each group, assemble the resulting group-wise BC e-values with appropriate weights, and feed the combined vector into the e-BH procedure. We show that the proposed method controls the FDR at the desired level in finite samples and achieves competitive power relative to state-of-the-art methods.
@@
Our approach involves a data-dependent method for weighting e-values when aggregating them from the BH and BC procedures (the second scenario) or assembling them from different subsets of the data (the first and third scenarios). It is important to note that our weighting method differs from the “boosting factor” proposed by Wang and Ramdas (2022), which involves multiplying each e-value by a factor to boost them up before applying the e-BH procedure. It is also worth mentioning the connection to the work of Ignatiadis et al. (2024b), where the authors used e-values as unnormalized weights for p-values in order to improve the testing power. The authors constructed these e-values using Basu’s theorem, which makes their weights independent of the p-values. In contrast, our weights are dependent on the e-values. The construction of our weights is motivated by the leave-one-out analysis for the BH and BC procedures. It ensures that the weighted e-values satisfy Condition (2) below, which is sufficient for the corresponding e-BH procedure to maintain FDR control at the desired level. Our numerical findings show that implementing the e-BH procedure with data-dependent weights improves its efficiency across all applications compared to the unweighted version.
@@
The remainder of the paper is organized as follows. Section 2 briefly reviews several multiple testing procedures and their relationship to the e-BH procedure. Sections 3-5 respectively present three distinct scenarios along with our newly proposed multiple testing methodologies tailored to different applications: (i) a multiple testing procedure controlling both group-wise and overall FDR, (ii) a hybrid knockoff method and a hybrid approach that integrates the BH and BC procedures, and (iii) a structure-adaptive multiple testing procedure. Section 6 provides concluding remarks. The Supplement includes additional numerical results as well as complete proofs of all main theoretical results.
@@ 2 Preliminaries
In recent work, Li and Zhang (2025) introduced a unified framework for understanding many commonly used multiple testing procedures and demonstrated their equivalence to the e-BH procedure using appropriately defined sets of e-values. In this section, we briefly review their results.
Consider $n$ hypotheses $H_{1},\dots,H_{n}$. Let $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$ denote the sets of true null and true alternative hypotheses, respectively. Let $\bm{\theta}=(\theta_{1},\dots,\theta_{n})\in\{0,1\}^{n}$ represent the true states of these hypotheses, where $\theta_{i}=0$ indicates that $H_{i}$ is under the null and $\theta_{i}=1$ otherwise. We define a decision rule $\bm{\delta}=(\delta_{1},\dots,\delta_{n})\in\{0,1\}^{n}$, where $\delta_{i}=1$ indicates rejection of $H_{i}$, and $\delta_{i}=0$ indicates acceptance.
The false discovery rate (FDR) associated with a decision rule $\bm{\delta}$ is defined as the expectation of the false discovery proportion (FDP):
@@
$\mathrm{FDR}(\bm{\delta})=\mathbb{E}[\mathrm{FDP}(\bm{\delta})],\quad\text{where}\quad\mathrm{FDP}(\bm{\delta})=\frac{\sum_{i=1}^{n}(1-\theta_{i})\delta_{i}}{1\vee\sum_{i=1}^{n}\delta_{i}},$
with $a\vee b=\max\{a,b\}$. An FDR-controlling procedure ensures that the FDR does not exceed a pre-specified threshold $\alpha\in(0,1)$.
2.1 Multiple Testing Procedures
Suppose we observe a set of p-values $p_{1},p_{2},\dots,p_{n}$ corresponding to the hypotheses $H_{1},H_{2},\dots,H_{n}$. Li and Zhang (2025) summarized several commonly used multiple testing procedures using the following unified form. Consider
$T=\sup\left\{t\in\mathcal{T}:\frac{m(t)}{1\vee\sum_{j=1}^{n}R_{j}(t)}\leq\alpha\right\},$ (1)
where $\mathcal{T}$ denotes the domain of the threshold, $m(t)$ provides a conservative estimate of the number of false rejections, and $R_{i}(t)$ indicates whether to reject the $i$th hypothesis at threshold $t$. The decision rule rejects hypothesis $H_{i}$ if and only if $R_{i}(T)=1$.
To specify an FDR-controlling procedure within this framework, one must define the functions $m(t)$ and $R_{i}(t)$. The most widely used method, the Benjamini-Hochberg (BH) procedure (Benjamini and Hochberg, 1995), sets $m(t)=nt$ and $R_{i}(t)=\mathbbm{1}\{p_{i}\leq t\}$, where $\mathbbm{1}\{A\}$ is the indicator function associated with set $A$. Storey’s (ST) procedure (Storey, 2002; Storey et al., 2004) refines the BH procedure by estimating the proportion of true null hypotheses from the observed p-values. Specifically, the ST procedure defines $R_{i}(t)=\mathbbm{1}\{p_{i}\leq t\}$ and $m(t)=n\pi_{0}^{\lambda}t$, where $\pi_{0}^{\lambda}\coloneqq\bigl{\{}1+n-R(\lambda)\bigr{\}}/\bigl{\{}(1-\lambda)n\bigr{\}}$ for a fixed $\lambda\in[0,1)$, and $R(\lambda)$ is the number of hypotheses rejected at threshold $\lambda$. Barber and Candès (2015) introduced the Barber-Candès (BC) procedure, a model-free approach leveraging symmetry properties of null p-values or test statistics to estimate false rejections. The BC procedure defines $m(t)=1+\sum_{i=1}^{n}\mathbbm{1}\{p_{i}\geq 1-t\}$ and $R_{i}(t)=\mathbbm{1}\{p_{i}\leq t\}$. The flexible BC (FBC) procedure proposed by Li and Zhang (2025) generalizes the BC approach using hypothesis-specific rejection functions $\varphi_{i}$, given by $m(t)=1+\sum_{i=1}^{n}\mathbbm{1}\{\varphi_{i}(1-p_{i})\leq t\}$ and $R_{i}(t)=\mathbbm{1}\{\varphi_{i}(p_{i})\leq t\}$. Table A.1 in Supplement A.1 summarizes the specifications of $m(t)$ and $R_{i}(t)$ for these procedures.
2.2 E-Values and the e-BH Procedure